1. A copper cube with sides of length 40mm had its surface blackened. The cube is heated up to 500K and place inside a room at a constant temperature of 300K. Calculate the initial rate of temperature fall of the cube, assuming it is a perfect blackbody. Take S.H.C of Aluminium to be 400J/kg/K, density of aluminium-7800kg/m'.


1
Expert's answer
2022-03-30T03:31:48-0400

"\frac Q {\Delta t S}=\frac{c\rho V\Delta T}{\Delta tS}=\sigma (T_1^4-T_2^4),"

"\frac{\Delta T}{\Delta t}=\frac{6\sigma (T_1^4-T_2^4)}{c\rho a}=0 .15~\frac Ks."